OpenStudy (anonymous):

find f`(x) for sqroot (x^(2)+4) some how i got x which i believe to be the wrong answer

OpenStudy (anonymous):

\[\sqrt{x ^{2}+4}\] here is what it looks like

OpenStudy (anonymous):

\[f'(x)=2x/(2(x^2+4))\]

OpenStudy (amistre64):

sqrt = ^1/2

OpenStudy (anonymous):

oh thats right

OpenStudy (anonymous):

so it would be x^2/2?

OpenStudy (amistre64):

\[(x^2+4)^{1/2}\] \[\frac{1}{2}(x^2+4)^{1/2-2/2}*(x^2+4)'\] \[\frac{1}{2}(x^2+4)^{1/2-2/2}*2x\]

OpenStudy (anonymous):

thank you both

OpenStudy (amistre64):

forgot to actually do the 1/2 - 2/2 lol

OpenStudy (amistre64):

\[\frac{1}{2}(x^2+4)^{1/2-2/2}*2x\] \[\frac{2}{2}x(x^2+4)^{-1/2}\] \[x(x^2+4)^{-1/2}=\frac{x}{\sqrt{x^2+4}}\]

OpenStudy (anonymous):

well i got x because i have horrible algebra skills :D

OpenStudy (anonymous):

so who know that these could be done a on the calculator